If you have 71 sweets and you give three sweets to each of your three friends, how many will you have left over?
Found 2 solutions by ewatrrr, heatherg2:
Answer by ewatrrr(24785)

71 - 3•3 = 71 - 9 = 62

Answer by heatherg2(1)

First of all, I'm going to figure out how many total treats that I'll be giving to my friends. If I give each one of my three friends three treats, that would be:
3 x 3 = 9 treats
Then I will just need to subtract 9 from the 71 treats that I started with to get how many are left.
71-9= 62 treats left over